sin-1 (x) = 2 tan-1(x). number of solutions are -


a) 1


b)2


c)3


d)5

Grade:12

1 Answers

man mohan mishra
12 years ago

   sin-1 x=2tan-1 x            or               sin-1 x=sin-1 (2x/1+x2 )             .                                       or                      x=2x/1+x2              .                                                            or               x+x3 =2x                    .                                        or              x3 -x=0                           .                                hence                      x=0,-1,1 are the solutions of the equation.                                                                                                    .     therefore, number of solutions are 3.